find the value of the discriminant and give the number of real solutions.
4(x^2+5x)=25

Answer by nabla(475)

4(x^2+5x)=25
4x^2+20x-25=0

The discriminant for a quadratic polynomial in form ax^2+bx+c is b^2-4ac.
You should be able to solve this problem for yourself at this point (and you should try), but the answer is as follows:
20^2-4(4)(-25)=400+400=800.